# EEG-SCMM: Soft Contrastive Masked Modeling for Cross-Corpus EEG-Based Emotion Recognition

### v1

@@ Discussion
Figure 5: Model performance with limited labeled data for fine-tuning on SEED and SEED-IV.
@@
We further explore the model performance of SCMM on the SEED and SEED-IV datasets when fine-tuning labeled data is limited. Specifically, we randomly select 1%, 5%, 10% and 20% of labeled samples from the fine-tuning dataset for model fine-tuning, while the remaining samples are used for testing. Figure 5 shows the classification accuracy curves. For the SEED-IV3 $\rightarrow$ SEED3 and SEED-IV4 $\rightarrow$ SEED3 experiments, SCMM achieves classification accuracies exceeding 70% with only 1% of labeled data. The accuracies significantly improve as the proportion of labeled samples increases, reaching close to 100% with 10% of labeled data. Meanwhile, our model achieves classification accuracies over 60% and 70% with only 1% of labeled data in the SEED3 $\rightarrow$ SEED-IV3 and SEED3 $\rightarrow$ SEED-IV4 experiments. The classification accuracies exceed 90% when fine-tuning with 10% of labeled samples in both experiments. In summary, experimental results indicate that SCMM maintains superior performance even with limited labeled data for fine-tuning, showing its outstanding robustness and potential in few-shot scenarios. The complete results and computational complexity analysis are provided in Appendix F.
Comparison of different masking strategies
Strategies SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
Random 90.30 / 08.80 84.63 / 10.99
Channel 90.25 / 08.68 85.91 / 10.97
Hybrid 91.61 / 07.56 87.24 / 08.35
Table 5: Comparative experiments of different masking strategies on the SEED and SEED-IV 3-category datasets.
To assess the impact of various masking strategies, we conduct comparative experiments on the SEED and SEED-IV 3-category datasets using three strategies: random, channel, and hybrid masking. Table 5 presents the experimental results, showing that the hybrid masking strategy achieves the highest accuracy and the lowest standard deviation in both experiments. This suggests that the integration of different masking approaches significantly improves the model performance and stability. More details on the different masking strategies are presented in Appendix D.
Comparison of soft CL in the original data space and embedding space
@@
While soft contrastive learning has been explored across various domains, most methods focus on computing soft assignments for contrastive loss in the embedding space (Dwibedi et al. 2021; Yèche et al. 2021). However, we argue that utilizing similarities in the original data space offers superior self-supervision and is particularly well-suited for EEG emotional data. To validate this, we conduct additional experiments on the SEED, SEED-IV and DEAP datasets to verify the effectiveness of soft contrastive learning in the original data space. Specifically, we modify the metric function $Dist(\cdot,\cdot)$ to use similarities between projected embeddings, shifting the computation of soft assignments from the original data space to the embedding space. Table 4 presents the experimental results, demonstrating that soft contrastive learning in the original data space consistently outperforms the embedding space in all experiments. Furthermore, this approach allows for pre-computing cosine similarities between original sample pairs offline, thus reducing computational resource requirements and improving training efficiency.
Visualization
Figure 6: (a) Heat maps of pairwise similarity matrices. (b) t-SNE visualization of the learned embeddings. Best viewed in color. Zoom in for better view.
@@
To evaluate whether sample-wise relationships are preserved in the encoder, we randomly select 100 test samples from the SEED dataset, and visualize the pairwise cosine similarity between sample pairs. Additionally, we select all test samples of one subject from the SEED dataset and visualize the learned embeddings of SCMM using t-SNE (Van der Maaten and Hinton 2008). Figure 6(a) presents heat maps of pairwise similarity matrices, where darker colors indicate higher similarity between samples. Traditional hard CL identifies only coarse-grained relationships across samples from different emotion categories, especially for the most challenging-to-recognize negative and neutral emotions. In contrast, SCMM effectively captures the fine-grained relationships between samples of different categories. Moreover, the results of t-SNE visualization in Fig. 6(b) indicate that the proposed SCMM better clusters samples within the same category and increases the inter-class distance compared to traditional hard CL, thus enhancing the classification performance. More visualization results can be found in Appendix G.
Conclusion
@@
This paper proposes a novel self-supervised pre-training framework, Soft Contrastive Masked Modeling (SCMM), for cross-corpus EEG-based emotion recognition. Unlike traditional contrastive learning models, SCMM integrates soft contrastive learning with a hybrid masking strategy to effectively capture the ”short-term continuity” characteristics inherent in human emotions, and produce stable and generalizable EEG representations. Additionally, an aggregator is developed to weightedly aggregate complementary information from multiple close samples, thereby enhancing fine-grained feature representation capability in the modeling process. Extensive experiments on three well-recognized datasets show that SCMM consistently achieves SOTA performance in cross-corpus EEG-based emotion recognition tasks under both same-class and different-class conditions. Comprehensive ablation study and parameter analysis confirm the superior performance and robustness of SCMM. Visualization results indicate that our model effectively reduces the distance between similar samples within the same category, and captures more fine-grained relationships across samples. These findings suggest that SCMM enhances the feasibility of extending the proposed method to real-life aBCI applications.
@@ Appendix A Datasets
A.1 Dataset Description
We conduct extensive experiments on three well-recognized datasets, SEED (Zheng and Lu 2015), SEED-IV (Zheng et al. 2018), and DEAP (Koelstra et al. 2011), to evaluate the model performance of SCMM in cross-corpus EEG-based emotion recognition tasks. Table 6 provides a detailed description of the three datasets.
@@ A.1 Dataset Description
SEED 15 3 $\times$ 15 62 1kHz 3 (Negative, Neutral, Positive)
SEED-IV 15 3 $\times$ 15 62 1kHz 4 (Sad, Neutral, Fear, Happy)
DEAP 32 1 $\times$ 40 32 512Hz Valence: 1 - 9, Arousal: 1 - 9
Table 6: Detailed description of the experimental datasets.
@@
(1) SEED (Zheng and Lu 2015) was developed by Shanghai Jiao Tong University. The dataset used a 62-channel ESI NeuroScan System based on the international 10-20 system to record EEG signals from 15 subjects (7 males and 8 females) under different video stimuli at a sampling rate of 1kHz. Each subject participated in 3 sessions. In each session, each subject was required to watch 15 movie clips consisting of 3 different emotional states: negative, neutral and positive. Each emotional state contains a total of 5 movie clips, corresponding to 5 trials.
@@
(2) SEED-IV (Zheng et al. 2018) used the same EEG acquisition equipment as the SEED dataset, but with different video stimuli, emotion categories and subjects. The dataset recorded EEG signals from 15 subjects under different video stimuli at a sampling rate of 1kHz. Each subject participated in 3 sessions. In each session, each subject was required to watch 24 movie clips containing 4 different emotions: sad, neutral, fear and happy. Each emotion contains a total of 6 movie clips, corresponding to 6 trials.
@@
(3) DEAP (Koelstra et al. 2011) was constructed by Queen Mary University of London. The dataset used a 128-channel Biosemi ActiveTwo System to record EEG signals from specific 32 channels of 32 subjects (16 males and 16 females) while watching 40 one-minute music videos at a sampling rate of 512Hz. The 40 one-minute videos elicited different emotions according to the valence-arousal dimension. Specifically, the valence-arousal emotional model, first proposed by Russell (Russell 1980), places each emotional state on a two-dimensional scale. The first dimension represents valence, ranging from negative to positive, and the second dimension represents arousal, ranging from calm to exciting. Participants rated valence and arousal using a continuous scale of 1 to 9 after watching each video clip.
A.2 Pre-processing Procedures
@@
For the SEED and SEED-IV datasets, the raw EEG signals were initially downsampled to 200Hz and filtered through a bandpass filter of 0.3-50Hz to filter noise and remove artifacts. Then, the data were divided into multiple non-overlapping segments using sliding windows of 1 second for SEED and 4 seconds for SEED-IV, respectively. After that, we extracted differential entropy (DE) features for each channel of each segment at five frequency bands: Delta (1-4Hz), Theta (4-8Hz), Alpha (8-14Hz), Beta (14-31Hz), and Gamma (31-50Hz). Finally, the DE features from 62 channels and 5 bands were formed into a feature matrix of shape $62\times 5$, which serves as input to the SCMM model. The extraction of DE features can be expressed as:
$DE(x)=\frac{1}{2}{\rm log}(2\pi e\sigma^{2}),$ (9)
@@
Here, $x$ represents an EEG signal segment of a specific length that approximately obeys a Gaussian distribution $N(\mu,\sigma^{2})$, where $\sigma$ denotes the standard deviation of $x$, and $e$ is the Euler constant.
@@
For the DEAP dataset, the raw EEG signals were first downsampled to 128Hz and denoised by a bandpass filter of 4-45Hz. Subsequently, the data were segmented into multiple non-overlapping segments using a sliding window of 1s. Similar to the SEED and SEED-IV datasets, DE features were extracted for each channel of each segment at five frequency bands. Finally, the DE features from 32 channels and 5 bands were formed into a feature matrix of shape $32\times 5$ as input to the model. During the experiments, we divided the continuous labels using a fixed threshold of 5 to convert them to binary classification tasks (low / high).
A.3 Handling Different Number of Channels
@@
Since the SEED-series datasets and the DEAP dataset contain different numbers of electrodes (channels), we require channel processing before inputting DE features into the model. Specifically, we consider the fine-tuning dataset as an anchor. When the number of channels in the fine-tuning dataset is less than in the pre-training dataset, we select data from the corresponding channels in the pre-training dataset and drop the data from the redundant channels as inputs (e.g., pre-training on SEED and fine-tuning on DEAP). Conversely, when the number of channels in the fine-tuning dataset is greater than in the pre-training dataset, we fill the missing channel data with zeros in the pre-training dataset to match the fine-tuning dataset (e.g., pre-training on DEAP and fine-tuning on SEED).
Appendix B Implementation Details
To reduce computational load while maintaining model performance, we adopt a lightweight design for each module of SCMM. Specifically, we use a 3-layer 1D CNN for the encoder $E$ and a 2-layer MLP for the projector $P$. For the lightweight decoder $D$, we utilize a single-layer MLP for reconstruction. Regarding hyperparameter selection in the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking, i.e., the percentage of random masking and channel masking is 0.9 and 0.1, respectively. We use the negative of cosine similarity as the metric function $Dist(\cdot,\cdot)$, and we set $\alpha$ to 0.5, $\tau_{w}$ to 0.05 and $\tau_{c}$ to 0.5 for soft contrastive learning. We use Adam as the optimizer with a learning rate of $5\times 10^{-4}$ and an L2-norm penalty coefficient $3\times 10^{-4}$. The pre-training process is conducted over 200 epochs with a batch size of 256. We save the model parameters $\theta$ from the final epoch as the pre-trained SCMM. In the fine-tuning stage, we input the encoded embeddings $h_{i}$ into a classifier consisting of a 2-layer fully connected network for final emotion recognition. The Adam optimizer is utilized with a learning rate of $5\times 10^{-4}$ and a weight decay of $3\times 10^{-4}$. The number of fine-tuning epochs is set to 50 for the SEED and SEED-IV datasets and 500 for the DEAP dataset, with a batch size of 128. For efficient deployment and testing of the model, the pre-trained SCMM is optimized solely using cross-entropy loss during fine-tuning. All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce 3090 GPU.
   Pre-training    Fine-tuning
@@ Appendix B Implementation Details
Table 7: Hyperparameter settings of SCMM.
Appendix C Baseline Methods and Experimental Settings
C.1 Baseline Methods
We compare the proposed SCMM against eight competitive methods, including four traditional deep learning methods: BiDANN (Li et al. 2018), TANN (Li et al. 2021), PR-PL (Zhou et al. 2023) and E2STN (Zhou et al. 2024), as well as four self-supervised learning models: SimCLR (Chen et al. 2020; Tang et al. 2021), Mixup (Zhang et al. 2018; Wickstrøm et al. 2022), MAE (He et al. 2022) and JCFA (Liu et al. 2024a). Details of the eight baseline methods are summarized as follows.
  • BiDANN (Li et al. 2018): The bi-hemispheres domain adversarial neural network mapped the EEG data of both left and right hemispheres into discriminative feature spaces separately to address domain shifts in EEG-based emotion recognition tasks.
  • TANN (Li et al. 2021): The transferable attention neutral network is a novel transfer learning methods which learned the discriminative information from EEG signals using local and global attention mechanisms.
  • PR-PL (Zhou et al. 2023): The prototypical representation based pairwise learning framework adopted pairwise learning to model the relative relationships between EEG sample pairs in terms of prototypical representations, addressing the critical issues of individual differences and noise labels in cross-subject scenarios.
  • E2STN (Zhou et al. 2024): The emotional EEG style transfer network integrated content information from the source domain with style information from the target domain, achieving superior performance in cross-corpus EEG-based emotion recognition tasks.
  • SimCLR (Chen et al. 2020; Tang et al. 2021): A seminal work in self-supervised contrastive learning, first proposed for computer vision, has been extended to human activity recognition.
  • Mixup (Zhang et al. 2018; Wickstrøm et al. 2022): A novel CL-based data augmentation method that aimed to correctly predict the mixing proportion of two samples, has been applied to time series analysis.
  • MAE (He et al. 2022): A groundbreaking work in the field of mask modeling, which proposed to reconstruct the masked portion based on the unmasked portion of the masked sample, has achieved remarkable success in a wide range of fields.
  • JCFA (Liu et al. 2024a): The joint contrastive learning framework performed joint contrastive learning across two domains to synchronize the time- and frequency-based embeddings of the same EEG sample in the latent time-frequency space, achieving state-of-the-art performance in cross-corpus scenarios.
To ensure a fair comparison, we adopt the same encoder, projector, decoder, and classifier structures for SimCLR, Mixup and MAE as used in SCMM. We use the default hyperparameters reported in the original paper for all methods in our experiments, unless otherwise specified. Additionally, for BiDANN, TANN, PR-PL, E2STN, MAE, and SCMM, the input samples are preprocessed 1-s DE features. In contrast, SimCLR, Mixup, and JCFA use preprocessed 1-s EEG signals as inputs, in accordance with the specific design of each model.
C.2 Experimental Settings
We adopt a cross-corpus subject-independent experimental protocol in the experiments, following the setup used in JCFA. Specifically, samples from one dataset are used for pre-training, while samples of each subject from another dataset are used individually for fine-tuning and testing. During the fine-tuning process, we employ a leave-trials-out setting, where samples from a part of trials in each session of each subject in the fine-tuning dataset are used for fine-tuning, and the remaining trials are used for testing. For example, SCMM is pre-trained on the SEED-IV dataset and fine-tuned on the SEED dataset. The allocation of samples for fine-tuning/testing is as follows: 9/6 for the SEED dataset, 12/6 for the SEED-IV 3-category dataset, 16/8 for the SEED-IV 4-category dataset, and 24/16 for the DEAP dataset. This approach effectively avoids information leakage. The detailed experimental settings and data division for pre-training and fine-tuning are summarized in Table 8.
@@ C.2 Experimental Settings
Same-Class SEED-IV3 $\rightarrow$ SEED3 SEED-IV, 3-category SEED: 9 / 6 trials in each session per subject
SEED3 $\rightarrow$ SEED-IV3 SEED SEED-IV, 3-category: 12 / 6 trials in each session per subject
Different-Class SEED-IV4 $\rightarrow$ SEED3 SEED-IV, 4-category SEED: 9 / 6 trials in each session per subject
SEED3 $\rightarrow$ SEED-IV4 SEED SEED-IV, 4-category: 16 / 8 trials in each session per subject
DEAP $\rightarrow$ SEED3 DEAP SEED: 9 / 6 trials in each session per subject
SEED3 $\rightarrow$ DEAP (Valence) SEED DEAP (Valence): 24 / 16 trials per subject
SEED3 $\rightarrow$ DEAP (Arousal) SEED DEAP (Arousal): 24 / 16 trials per subject
Table 8: Pre-training and fine-tuning scenarios under two conditions for cross-corpus EEG-based emotion recognition.

### v2

@@ 6.1. Ablation Study
Figure 5. Model performance with limited labeled data for fine-tuning on SEED and SEED-IV under same-class and different-class cross-corpus conditions. Best viewed in color. Zoom in for a better view.
[…]
Model performance with limited labeled data for fine-tuning on SEED and SEED-IV.
6.2. Model Performance with Limited Data
@@
We investigate the model performance of SCMM on the SEED and SEED-IV datasets when fine-tuning with limited labeled data. Specifically, we randomly select 1%, 5%, 10%, and 20% of labeled samples from the fine-tuning dataset for model fine-tuning, while the remaining samples are used for testing. Figure 5 shows the classification accuracy curves. For the SEED-IV3 $\rightarrow$ SEED3 and SEED-IV4 $\rightarrow$ SEED3 experiments, SCMM achieves classification accuracies exceeding 70% with only 1% of labeled data. The accuracies significantly improve as the proportion of labeled samples increases, reaching close to 100% with 10% of labeled data. Meanwhile, our model achieves classification accuracies over 60% and 70% with only 1% of labeled data in the SEED3 $\rightarrow$ SEED-IV3 and SEED3 $\rightarrow$ SEED-IV4 experiments. The accuracies exceed 90% when fine-tuning with 10% of labeled samples in both experiments. In summary, the results indicate that SCMM maintains superior performance even with limited labeled data for fine-tuning, showing its outstanding robustness and potential in few-shot scenarios.
Figure 6. Hyperparameter sensitivity analysis. (a) - (f) represent the masking ratio $r$, threshold $\mu$, metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{\rm s}$, and temperature $\tau_{\rm c}$, respectively. Best viewed in color.
[…]
Sensitivity analysis of hyperparameters.
6.3. Delve into Contrastive Learning
6.3.1. Hard CL vs. Soft CL
We compare the model performance of SCMM using traditional hard CL and soft CL on SEED, SEED-IV, and DEAP under same-class and different-class cross-corpus conditions to verify the effectiveness of the proposed soft weighting mechanism. Table 4 presents the comparison results, showing that the SCMM model using the soft weighting mechanism consistently achieves the highest classification accuracies and lowest standard deviations in all experiments. This improvement indicates that assigning soft weights (rather than hard values) to different sample pairs when computing the contrastive loss effectively captures the inter-sample similarity relationships, thereby improving the discriminative ability of the model.
6.3.2. Soft CL in the Embedding Space vs. Original Data Space
@@ 6.3. Delve into Contrastive Learning
While soft CL has been explored across various domains, most methods focus on computing soft assignments for contrastive loss in the embedding space (Dwibedi et al., 2021; Yèche et al., 2021). However, we argue that utilizing similarities in the original data space can provide better self-supervision and is particularly suitable for emotional EEG data. To validate this, we conduct experiments on SEED, SEED-IV, and DEAP under same-class and different-class cross-corpus conditions to compare soft CL in different spaces. Specifically, we modify the metric function $Dist(\cdot,\cdot)$ to use similarities between projected embeddings, shifting the computation of soft assignments from the original data space to the embedding space. Table 5 presents the experimental results, demonstrating that soft CL in the original data space consistently outperforms the embedding space in all experiments. Furthermore, this approach allows offline pre-computation of cosine similarities of different sample pairs in the original data space, thus reducing computational costs and improving training efficiency.
6.4. Hyperparameter Sensitivity Analysis
To analyze the hyperparameter sensitivity of SCMM, we conduct experiments on SEED and SEED-IV under the same-class condition, as shown in Fig. 6. Specifically, the examined hyperparameters are divided into two groups: (1) Hybrid Masking: masking ratio $r$ and threshold $\mu$, and (2) Soft Contrastive Learning: metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{\rm s}$, and temperature $\tau_{\rm c}$.
(1) Hybrid Masking. Figures 6(a) - (b) show the classification curves of SCMM with different masking ratios $r$ and thresholds $\mu$. Experimental results show that the proposed SCMM achieves the best performance when the masking ratio and threshold are set to $r=0.5$ and $\mu=0.1$ (i.e., the ratio of random masking and channel masking is 9:1), respectively. In addition, we find that our model performs better when using hybrid masking compared to random masking or channel masking in most settings, demonstrating the effectiveness of the proposed hybrid masking strategy in enhancing the model performance.
(2) Soft Contrastive Learning. Figures 6(c) - (f) depict the classification curves of SCMM using different metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{\rm s}$, and temperature $\tau_{\rm c}$, respectively. Specifically, our model performs best on the SEED and SEED-IV 3-category datasets when using cosine similarity as the metric function $Dist(\cdot,\cdot)$, as shown in Fig. 6(c). Additionally, Figures 6(d) - (f) show that SCMM achieves the best performance when the upper bound, sharpness and temperature are set to $\mu=0.5$, $\tau_{\rm s}=0.05$ and $\tau_{\rm c}=0.5$, respectively. In summary, the results demonstrate that SCMM maintains excellent performance under different hyperparameter settings, indicating that its superior generalization ability is not significantly affected by hyperparameter changes.
6.5. Visualization
6.5.1. Sample-Wise Relationships
@@
To evaluate whether sample-wise relationships are preserved in the encoder, we randomly select 100 test samples from the SEED dataset and visualize the pairwise cosine similarity between sample pairs. Additionally, we select all test samples of one subject from the SEED dataset and visualize the learned embeddings of SCMM using t-SNE (Van der Maaten and Hinton, 2008). Figure 7(a) presents heat maps of pairwise similarity matrices, where darker colors indicate higher similarity between samples. Traditional hard CL identifies only coarse-grained relationships across samples from different emotion categories, especially for the most challenging-to-recognize negative and neutral emotions. In contrast, SCMM effectively captures the fine-grained relationships between samples of different categories. Moreover, the results of t-SNE visualization in Fig. 7(b) indicate that our model better clusters samples within the same category and increases the inter-class distance compared to hard CL, thus enhancing the classification performance.
Figure 7. (a) Heat maps of pairwise similarity matrices. (b) t-SNE visualization of the learned embeddings. Best viewed in color. Zoom in for a better view.
[…]
Visualization.
6.5.2. Intra- and Inter-Class Similarities
Figure 8. Intra- and inter-class cosine similarities of embeddings learned by hard CL and soft CL. (a) and (b) represent the average intra-class and inter-class cosine similarities, respectively. Best viewed in color.
[…]
Intra- and inter-class cosine similarities.
Figure 9. Comparison of reconstruction quality. We visualize the reconstruction results to compare the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM). Best viewed in color.
[…]
Comparison of reconstruction quality.
To assess the quality of the embeddings learned by SCMM, we randomly select one subject from the SEED dataset and calculate both the average intra- and inter-class cosine similarities between the learned embeddings of all test samples, as shown in Fig. 8. It is evident that the proposed SCMM produces embeddings with higher intra-class similarity compared to traditional hard CL. In addition, the average inter-class similarity of the embeddings learned by SCMM is significantly lower than that of hard CL. In summary, visualization results confirm that the soft contrastive learning strategy designed in SCMM effectively clusters samples within the same category while distinctly separating samples from different categories, thus enhancing the model’s discriminative capabilities.
6.5.3. Reconstruction Quality
To verify the effectiveness of the similarity-aware aggregator designed in SCMM, we compare the reconstruction quality of the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM) on the DEAP dataset. For clarity, we flatten both the original input sample and the reconstructed sample into one-dimensional vectors with dimensions $C\times F$ (channels $\times$ features). The results depicted in Fig. 9 illustrate that our model achieves lower reconstruction loss (MSE) and better sample reconstruction, which is crucial to improving the model performance.
7. Conclusions
@@
This paper proposes a novel SCMM framework to tackle the critical challenge of cross-corpus generalization in the field of EEG-based emotion recognition. Unlike traditional CL-based models, SCMM integrates soft contrastive learning with a hybrid masking strategy to effectively capture the ”short-term continuity” characteristics inherent in human emotions and produce stable and generalizable EEG representations. Additionally, a similarity-aware aggregator is introduced to fuse complementary information from semantically related samples, thereby enhancing the fine-grained feature representation capability of the model. Extensive experimental results on three well-recognized datasets show that SCMM consistently achieves SOTA performance in cross-corpus EEG-based emotion recognition tasks under both same-class and different-class conditions. Comprehensive ablation study and hyperparameter sensitivity analysis confirm the superiority and robustness of SCMM. Visualization results indicate that our model effectively reduces the distance between similar samples within the same category and captures more fine-grained relationships across samples. These findings suggest that SCMM enhances the feasibility of extending the proposed method to real-life aBCI applications.
8. Acknowledgements
This work was funded by the National Natural Science Foundation of China (62276169), the STI 2030-Major Projects (2021ZD0200500), the Medical-Engineering Interdisciplinary Research Foundation of Shenzhen University (2023YG004), the Shenzhen-Hong Kong Institute of Brain Science-Shenzhen Fundamental Research Institutions (2023SHIBS0003), the Open Research Fund of the State Key Laboratory of Brain-Machine Intelligence, Zhejiang University (Grant No. BMI2400008), and the Shenzhen Science and Technology Program (No. JCYJ20241202124222027 and JCYJ20241202124209011).
@@ Appendix A Datasets
A.1. Dataset Description
We conduct extensive experiments on three well-known datasets, SEED (Zheng and Lu, 2015), SEED-IV (Zheng et al., 2018), and DEAP (Koelstra et al., 2011), to evaluate the model performance of SCMM in cross-corpus EEG-based emotion recognition tasks. These datasets encompass different EEG equipment, emotional stimuli, data specifications, labeling approaches, and subjects, making them exceptionally suitable for evaluating the model’s efficacy in cross-corpus scenarios. Table 6 provides a detailed description of the three datasets.
Table 6. Detailed descriptions of the experimental datasets.
@@ A.1. Dataset Description
SEED (Zheng and Lu, 2015) 15 3 $\times$ 15 62 1 kHz 3 (Negative, Neutral, Positive)
SEED-IV (Zheng et al., 2018) 15 3 $\times$ 24 62 1 kHz 4 (Sad, Neutral, Fear, Happy)
DEAP (Koelstra et al., 2011) 32 1 $\times$ 40 32 512 Hz Valence: 1 - 9, Arousal: 1 - 9
@@
(1) SEED (Zheng and Lu, 2015) was developed by the Center for Brain-like Computing and Machine Intelligence (BCMI) of Shanghai Jiao Tong University. The dataset used a 62-channel ESI NeuroScan System based on the international 10-20 system to record EEG signals from 15 subjects (7 males and 8 females) under different video stimuli at a sampling rate of 1 kHz. Each subject participated in 3 sessions. In each session, each subject was required to watch 15 movie clips consisting of 3 different emotional states: negative, neutral, and positive. Each emotional state contains a total of 5 movie clips, corresponding to 5 trials.
@@
(2) SEED-IV (Zheng et al., 2018) used the same EEG equipment as the SEED dataset, but with different video stimuli, emotion categories, and subjects. The dataset recorded EEG signals from 15 subjects under different video stimuli at a sampling rate of 1 kHz. Each subject participated in 3 sessions. In each session, each subject was required to watch 24 movie clips containing 4 different emotions: sad, neutral, fear, and happy. Each emotion contains a total of 6 movie clips, corresponding to 6 trials.
@@
(3) DEAP (Koelstra et al., 2011) was constructed by Queen Mary University of London. The dataset has completely different acquisition device, emotional stimuli, data specifications, labeling approaches, experimental protocols, and subjects from the SEED-series datasets. Specifically, the DEAP dataset used a 128-channel Biosemi ActiveTwo System to record EEG signals from specific 32 channels of 32 subjects (16 males and 16 females) while watching 40 one-minute music videos at a sampling rate of 512 Hz. The 40 videos elicited different emotions according to the valence-arousal dimension. Specifically, the valence-arousal emotional model proposed by Russell (Russell, 1980) places each emotional state on a two-dimensional scale. The first dimension represents valence, ranging from negative to positive, and the second dimension represents arousal, ranging from calm to exciting. Participants rated valence and arousal using a continuous scale of 1 to 9 after watching each video clip.
A.2. Pre-processing Procedures
@@
For the SEED and SEED-IV datasets, the raw EEG signals were first downsampled to 200 Hz and filtered through a bandpass filter of 0.3-50 Hz to remove noise and artifacts. Then, the data were divided into multiple non-overlapping segments using a sliding window of 1s (SEED) and 4s (SEED-IV), respectively. After that, we extracted differential entropy (DE) features for each channel of each segment at five frequency bands: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz). Finally, the DE features from 62 channels and 5 bands were formed into a feature matrix of shape 62×562$\times$ 562 × 5, which serves as input to the SCMM model. The extraction of DE features can be expressed as:
(9) $DE(x)=\frac{1}{2}{\rm log}(2\pi e\sigma^{2}),$
@@
where $e$ is the Euler constant. $x$ is an EEG signal of a specific length that approximately obeys a Gaussian distribution $N(\mu,\sigma^{2})$. Here, $\mu$ and $\sigma$ are the mean and standard deviation of $x$, respectively.
@@
For the DEAP dataset, the raw EEG signals were initially downsampled to 128 Hz and denoised by a bandpass filter of 4-45 Hz. Subsequently, the data were segmented into multiple non-overlapping segments using a sliding window of 1s. Similar to the SEED and SEED-IV datasets, DE features were extracted for each channel of each segment at five frequency bands. Finally, the DE features from 32 channels and 5 bands were formed into a feature matrix of shape 32×532$\times$ 532 × 5 as input to the model. In the experiments, we divided the continuous labels using a fixed threshold of 5 to convert them to binary classification tasks (low/high).
A.3. Handling Different Number of Channels
@@
Since the SEED-series datasets and the DEAP dataset contain different numbers of electrodes (channels), we require channel processing before inputting DE features into the model. Specifically, we consider the fine-tuning dataset as the standard. When the number of channels in the fine-tuning dataset is less than in the pre-training dataset, we select data from the corresponding channels in the pre-training dataset and drop the data from the redundant channels as inputs (e.g., pre-training on SEED and fine-tuning on DEAP). Conversely, when the number of channels in the fine-tuning dataset is greater than in the pre-training dataset, we fill the missing channel data with zeros in the pre-training dataset to match the fine-tuning dataset (e.g., pre-training on DEAP and fine-tuning on SEED).
Table 7. Hyperparameter settings of SCMM.
Hyperparameters Pre-training Fine-tuning
@@
Table 8. Experimental scenarios and data division for cross-corpus EEG-based emotion recognition.
@@
Same-Class SEED-IV3 $\rightarrow$ SEED3 SEED-IV, 3-class SEED: 9/6 trials in each session of each subject
SEED3 $\rightarrow$ SEED-IV3 SEED SEED-IV, 3-class: 12/6 trials in each session of each subject
Different-Class SEED-IV4 $\rightarrow$ SEED3 SEED-IV, 4-class SEED: 9/6 trials in each session of each subject
SEED3 $\rightarrow$ SEED-IV4 SEED SEED-IV, 4-class: 16/8 trials in each session of each subject
DEAP $\rightarrow$ SEED3 DEAP SEED: 9/6 trials in each session of each subject
SEED3 $\rightarrow$ DEAP (Valence) SEED DEAP (Valence): 24/16 trials of each subject
SEED3 $\rightarrow$ DEAP (Arousal) SEED DEAP (Arousal): 24/16 trials of each subject
Appendix B Implementation Details
To reduce computational load while maintaining model performance, we adopt a lightweight design for each module of SCMM. Specifically, we use a 3-layer 1D CNN for the encoder $E$ and a 2-layer MLP for the projector $P$. For the lightweight decoder $D$, we utilize a single-layer MLP for reconstruction. For the hyperparameter selection in the pre-training stage, we set $r$ to 0.5 and $\mu$ to 0.1 for hybrid masking, i.e., the ratio of random masking and channel masking is 9:1. We use the negative of cosine similarity as the metric function $Dist(\cdot,\cdot)$, and we set $\alpha$ to 0.5, $\tau_{w}$ to 0.05 and $\tau_{c}$ to 0.5 for soft contrastive learning. We use Adam optimizer with an initial learning rate of $5\times 10^{-4}$ and an L2-norm penalty coefficient $3\times 10^{-4}$. The pre-training process is conducted over 200 epochs with a batch size of 256. We save the model parameters $\theta$ from the final epoch as the pre-trained SCMM. In the fine-tuning stage, we input the encoded embeddings $h_{i}$ into an emotion classifier consisting of a 2-layer MLP for final emotion recognition. The Adam optimizer is utilized with an initial learning rate of $5\times 10^{-4}$ and a weight decay of $3\times 10^{-4}$. The number of fine-tuning epochs is set to 50 for SEED and SEED-IV and 500 for DEAP, with a batch size of 128. For efficient deployment and testing of the model, the pre-trained SCMM is optimized solely using cross-entropy loss during fine-tuning. All experiments are conducted using Python 3.9 with PyTorch 1.13 on an NVIDIA GeForce 3090 GPU. Table 7 summarizes the hyperparameter settings of SCMM.
Appendix C Baselines and Experimental Setup
We compare the proposed SCMM against ten competitive baselines, including five transfer learning methods: DANN (Ganin et al., 2016), BiDANN (Li et al., 2018b), TANN (Li et al., 2021), PR-PL (Zhou et al., 2023), and E2STN (Zhou et al., 2025), as well as five self-supervised learning models: SimCLR (Chen et al., 2020; Tang et al., 2020), Mixup (Zhang, 2017; Wickstrøm et al., 2022), TS-TCC (Eldele et al., 2021), MAE (He et al., 2022a), and JCFA (Liu et al., 2024). Note that E2STN and JCFA are two state-of-the-art (SOTA) methods designed for cross-corpus EEG-based emotion recognition. To ensure a fair comparison, we adopt the same encoder, projector, decoder, and classifier structures as SCMM for SimCLR, Mixup, TS-TCC, and MAE. We use the default hyperparameters reported in the original papers for all models in the experiments, unless otherwise specified. Additionally, for DANN, BiDANN, TANN, PR-PL, E2STN, MAE, and SCMM, the inputs are preprocessed 1-s DE features. In contrast, SimCLR, Mixup, TS-TCC, and JCFA use preprocessed 1-s EEG signals as inputs, in accordance with the specific design of each model.
[…]
Figure 10. Examples of generated masked samples using different masking strategies. The masking ratio is set to $r=0.5$, and the thresholds for hybrid masking are set to $\mu=0.1,0.5$, and 0.90.90.9, respectively. Best viewed in color.
[…]
Examples of generated masked samples using different masking strategies.
Table 9. Comparison of different masking strategies on SEED and SEED-IV under the same-class condition.
Strategies SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
Random 90.30 / 08.80 84.63 / 10.99
Channel 90.25 / 08.68 85.91 / 10.97
Parallel 90.34 / 08.20 86.23 / 09.94
Hybrid 91.61 / 07.56 87.24 / 08.35